Amy has three older sisters. Beth is 2 years older than Amy. Jo is 3 years older than Beth. Meg is 4 years older than Jo. Meg is twice as old as Beth. How old is Amy?
Answer by mananth(16949)

Amy has three older sisters. Beth is 2 years older than Amy. Jo is 3 years older than Beth. Meg is 4 years older than Jo. Meg is twice as old as Beth. How old is Amy?
Amy = x
Beth = x+2
Jo = x+2+3==>x+5
Meg = x+5+4===> = x+9
...
(x+9) = 2(x+2)
x+9= 2x+4
x=5 Amy's age